OpenStudy (anonymous):

how do you find x squared-8x+16 ?? help me please!

OpenStudy (hoblos):

= (x-4)^2

hero (hero):

x^2-8x+16 = 0 x^2-4x-4x + 16 = 0 x(x-4)-4(x-4) = 0 (x-4)(x-4) = 0 (x-4)^2 = 0 x-4 = 0 x = 4

OpenStudy (anonymous):

Thank you! but i dont need to solve for x. I have to factor the quadratic expressions

hero (hero):

You should post your questions more clearly. Seemed like you wanted to find x.

OpenStudy (anonymous):

oh ok next time I will. well I need to factor quadratic expressions. so how do I factor x squared - 8x + 16 ??

hero (hero):

I posted the steps above

OpenStudy (anonymous):

ohhhhhhhhh.... thanks

hero (hero):

Where's my medal?

OpenStudy (anonymous):

there